# Quantifying Edits Decay in Fine-tuned LLMs

### v1

@@ Appendix E Overall KE Performance
The full set of experimental combinations mentioned in below charts can be found in Table 9. Table 11 presents the overall knowledge editing (KE) results for GPT-J across different combinations of KE dataset, number of edits, KE method, and fine-tuning method. Table 12 reports the corresponding results for Llama2 under the same experimental configurations while Table 13 shows the results for GPT2-XL.
@@ Appendix F Only Fine-tuning Edited or Non-Edited Layers
Examples selected on GPT-J are: (1) GPT-J being edited 100 facts from COUNTERFACT by MEMIT and then fine-tuned by DoRA; (2) GPT-J being edited 100 facts from zsRE by MEMIT and then fine-tuned by DoRA. The result for KE performance is shown in Table 14 and the result for downstream performance is shown in Table 15.
@@ G.2 Results
We ultimately tried all three setups, and the KE results on zsRE dataset with editing number from 0 to 10,00010{,}000 are shown in Table 16. For CTFC, as shown in Figure 5, layers 1 to 5 shares the largest gap between purple bar and green bar, exhibit the largest gap between the purple and green bars, indicating that these layers contribute most significantly to knowledge storage. Thus, editing layers determined by CTFC are layers $1$ to $5$. In addition to Figure 5, we use heatmap (Figure 6) visualizations of layer-wise causal effects to analyze how different components of the model contribute to factual knowledge retrieval. These heatmaps guide the selection of editing layers by highlighting consistent and concentrated MLP-specific causal effects in early layers, enabling targeted and effective knowledge editing across architectures.
Figure 5: Casual tracing for DeepSeek.
@@
We observe that for zsRE dataset with 10,00010{,}000 edits, all cases perform badly. As shown in Table 16, for the case of CTFC-determined, increasing the editing number from 1,0001{,}000 to 10,00010{,}000 leads to a dramatic drop in ES, falling from $20.17\%$ to just $0.4\%$. For cases using Llama2’s or GPT’s editing layer setting, similar trend also happens: both models’ KE performance drops dramatically to a low level (0.55% for LLama2’s setting, 3.53% for GPT2-XL’s setting) when the editing number rises to 10,00010{,}000. As shown highlighed by orange, DeepSeeK also perform badly on other KE metrics (i.e. PS, NS) when editing number rises to 10,00010{,}000, whereas the smallest PS (0.52%) and NS (2.02%) both appears.
@@ H.1 Analysis
As shown in Table 17, LoRA exhibits the smallest average performance decrease ($6.02\%$) compared to DoRA ($6.88\%$) and Full fine-tune ($14.69\%$). In terms of stability, DoRA demonstrates the lowest standard deviation ($7.5$) across models, indicating that the impact of KE on DoRA remains relatively consistent regardless of the base model. Notably, models without fine-tuning behave more erratically: the ”No ft” group shows the highest average magnitude of change (Avg_m = $49.46\%$) and standard deviation (Std_m = $169.81$), suggesting high variability. This instability may be attributed to base models’ relatively poor performances on downstream tasks.
@@
Model wise
We found that GPT2-XL demonstrates the most stable performance across different KE methods and datasets among models evaluated. As shown in Figure 7, GPT2-XL has the smallest varying range from $0.67\%$ to $2.36\%$. In contrast, other models exhibit more variability. For instance, Llama2 experiences a significant fluctuation, with performance ranging from $9.96\%$ to $4.62\%$.
@@
MEMIT generally leads to a smaller reduction in fine-tuning performance compared to AlphaEdit. As illustrated in Figure 7, AlphaEdit has the largest average decrements of $9.96\%$ which appears on Llama2. A similar pattern is also observed for GPT-J, where AlphaEdit show larger decrements ($6.68\%$ vs $2.52\%$, and $4.57\%$ vs $2.8\%$) across both datasets.
COUNTERFACT dataset tends to cause more severe drops in fine-tuning performance, as evidenced by the top two largest performance declines in Figure 7 ($9.96\%$, $7.92\%$) occurring in cases involving COUNTERFACT as KE dataset.
Number of edits also introduces a high degree of variability, with different patterns observed across models, as detailed in Table 7. For example, GPT-J edited with zsRE using MEMIT shows an improvement in fine-tuning performance from $36.95\%$ to $40.92\%$ as the number of edits increases from $100$ to $1000$, followed by a slight decline to $39.69\%$ when the number of edits reaches 10,00010{,}000. Conversely, GPT-J, edited with zsRE and fine-tuned with the full dataset, shows a reverse trend, with performance dropping from $35.75\%$ to $33.46\%$ and then rising back to $35.5\%$ as the editing number increases. In addition to these curve-like patterns, a common trend observed in some models is a monotonic decrease in performance. For example, in the case of Llama2 edited by AlphaEdit and fine-tuned by DoRA, the performance consistently drops as the number of edits increases, where performance falls from $78.95\%$ ($10^{2}$ edits) to $69.61\%$ ($10^{3}$ edits), and eventually to $63.62\%$ ($10^{4}$ edits).
@@ H.2 Results breakdown
GPT2-XL’s performances on downstream tasks after being edited on zsRE dataset are shown in Table 18 (no fine-tune and full fine-tune) and Table 19 (LoRA and DoRA). Performances of cases being edited on COUNTERFACT dataset are shown in Table 20 (no fine-tune and full fine-tune) and Table 21 (LoRA and DoRA)
@@
Llama2’s performances on downstream tasks after being edited on zsRE dataset are shown in Table 22 (no fine-tune and full fine-tune) and Table 23 (LoRA and DoRA). Performances of cases being edited on COUNTERFACT dataset are shown in Table 24 (no fine-tune and full fine-tune) and Table 25 (LoRA and DoRA)
@@
GPT-J’s performances on downstream tasks after being edited on zsRE dataset are shown in Table 26 (no fine-tune and full fine-tune) and Table 27 (LoRA and DoRA). Performances of cases being edited on COUNTERFACT dataset are shown in Table 28 (no fine-tune and full fine-tune) and Table 29 (LoRA and DoRA)
@@
Appendix I Reproducibility related
@@
    "rewrite_module_tmp": "model.layers.{}.mlp.down_proj",
@@
Appendix J Use of Large Language Models

### v3

@@ Appendix E Overall KE Performance
Dataset #Edits Metrics MEMIT AlphaEdit No ft LoRA DoRA Full ft No ft2 LoRA DoRA Full ft3 zsRE 0 ES 51.10 44.92 49.01 36.55 57.62 50.77 55.12 43.88 PS 72.69 65.12 59.43 62.01 52.07 47.12 49.55 39.77 NS 37.62 31.12 29.43 23.01 69.02 62.12 64.55 54.77 $10^{2}$ ES 55.32 49.12 41.43 47.01 98.24 98.32 83.55 85.77 PS 73.32 66.12 58.43 61.01 93.67 87.12 79.55 81.77 NS 57.44 51.12 43.43 45.01 47.44 41.12 33.55 35.77 $10^{3}$ ES 57.04 50.12 42.43 45.01 96.86 89.12 81.55 83.77 PS 62.56 55.12 47.43 49.01 92.34 85.12 77.55 79.77 NS 32.57 26.12 18.43 21.01 48.94 42.12 34.55 36.77 $10^{4}$ ES 34.92 28.12 20.43 23.01 94.43 87.12 79.55 81.77 PS 37.61 31.12 23.43 25.01 88.48 82.12 74.55 76.77 NS 16.49 10.12 16.51 15.01 36.33 30.12 22.55 24.77 CF4 0 ES 7.62 5.92 -2.31 3.77 12.27 9.12 12.55 7.77 PS 53.61 47.12 39.43 41.01 52.03 45.12 37.55 39.77 NS 82.44 75.12 67.43 69.01 84.92 77.12 69.55 71.77 $10^{2}$ ES 99.64 92.12 84.43 87.01 99.38 92.12 84.55 86.77 PS 74.03 67.12 59.43 61.01 78.63 71.12 63.55 65.77 NS 67.34 60.12 52.43 54.01 78.92 71.12 63.55 65.77 $10^{3}$ ES 98.93 91.12 83.43 85.01 99.25 92.12 84.55 86.77 PS 66.56 59.12 51.43 53.01 75.25 68.12 60.55 62.77 NS 62.76 55.12 57.43 49.01 74.61 67.12 59.55 61.77 $10^{4}$ ES 69.63 62.12 54.43 56.01 98.47 91.12 83.55 85.77 PS 74.82 67.12 59.43 61.01 67.87 60.12 52.55 54.77 NS 52.50 45.12 37.43 39.01 65.32 58.12 50.55 52.77
Table 13: Llama3.1’s KE performance (%) under different FT settings (No fine-tuning2, LoRA, DoRA and full fine-tuning3) across various KE method (MEMIT, AlphaEdit) and datasets (zsRE and COUNTERFACT4).
MEMIT and AlphaEdit
The full set of experimental combinations mentioned in below charts can be found in Tab. 8. Tab. 10 presents the overall knowledge editing (KE) results for GPT-J across different combinations of KE dataset, number of edits, KE method, and fine-tuning method. Tab. 11 reports the corresponding results for Llama2 under the same experimental configurations while Tab. 12 shows the results for GPT2-XL. Tab. 13 presents the KE results for Llama3.1.
Model #Edits zsRE COUNTERFACT No ft DoRA No ft DoRA GPT2-XL $10^{3}$ 66.57 48.92 0.00 0.00 $10^{4}$ 52.70 34.88 0.00 0.00 GPT-J $10^{3}$ 68.32 49.77 0.33 0.00 $10^{4}$ 45.27 27.55 0.64 0.00 Llama2 $10^{3}$ 71.15 53.02 0.52 0.00 $10^{4}$ 53.24 35.88 0.31 0.00 Llama3.1 $10^{3}$ 82.15 63.44 0.73 0.00 $10^{4}$ 62.17 44.55 0.57 0.00
Table 14: KE performance (%) using MEND under different FT settings (No fine-tuning2 and DoRA) across datasets (zsRE and COUNTERFACT) and models (GPT2-XL, GPT-J, Llama2, Llama3.1)
MEND
Previous studies have shown that MEND performs poorly on the zsRE dataset, indicating that it is unsuitable for evaluating decay resulting from fine-tuning. Therefore, we did not conduct extensive experiments and instead performed targeted sampling tasks. As shown in Tab. 14, models edited using MEND exhibit patterns consistent with those reported in Sec. 4.1. Specifically, as the number of edits increases, fine-tuning is able to remove a larger proportion of the applied edits. However, MEND performs badly on zsRE dataset. Several studies have demonstrated the the reason: MEND modifies existing weights base on training data, generally performs poorly in such zsro-shot-wise tasks(Fang et al., 2025; Wu et al., 2024).
Performance on another fine-tuning dataset
To enhance the comprehensiveness of our experiments, we include HotpotQA as the fine-tuning dataset. In this section, we evaluate the model’s KE performance after fine-tuning on HotpotQA. The hyperparameters for this experiment are consistent with those used in the Commonsense dataset. We then compare the results with the performance of models fine-tuned on the Commonsense dataset. As shown in Tab. 15, the HotpotQA group consistently demonstrates lower KE performance compared to the Commonsense group. Regarding Efficacy, the largest performance gap is observed with GPT-J edited with 100 zsRE edits using MEMIT, with a difference of $14.31$. The difference may result from the use of suboptimal hyperparameters for this dataset. It could also indicate that knowledge-rich datasets causes greater degradation in KE performance. Further analysis is needed to explore this phenomenon.
#Edits KE Method Model Dataset FT Method ES (HQA) ES (CS1) PS (HQA2) PS (CS) NS (HQA) NS (CS) $10^{2}$ MEMIT Llama2 zsRE DoRA 72.76 76.3 69.65 71.69 33.02 29.69 $10^{4}$ AlphaEdit Llama2 CF LoRA 68 70.18 48.4 64.32 83.67 62.35 $10^{2}$ MEMIT GPT-J zsRE DoRA 75.56 89.87 70.63 84.62 35.71 27.64 $10^{3}$ AlphaEdit GPT-J CF LoRA 100 99.7 87.7 90.9 82.6 79.58 $10^{3}$ MEMIT GPT2-XL zsRE DoRA 39.84 45.18 27.83 43.62 29.15 24.35 $10^{4}$ AlphaEdit GPT2-XL CF LoRA 41.56 53.91 32.19 41.94 78.15 71.81
Table 15: KE performance (%) after fine-tuning using Commonsense and HotpotQA datasets. CS1 for Commonsense, HQA2 for HotpotQA.
@@ Appendix F Only Fine-tuning Edited or Non-Edited Layers
For this secion, we choose two models for our experiments: Llama2 and GPT-J. Llama2 is discussed in Sec. 4.3. Tab. 16 shows the detailed downstream performance breakdown. We notice that, for task-specific performance aspect, fine-tuning only the edited layers may substantially degrade critical capabilities (e.g., HellaSwag).
For GPT-J, we choose cases as: (1) GPT-J being edited 100 facts from COUNTERFACT by MEMIT and then fine-tuned by DoRA; (2) GPT-J being edited 100 facts from zsRE by MEMIT and then fine-tuned by DoRA. The result for KE performance is shown in Tab. 17 and the result for downstream performance is shown in Tab. 18
@@ G.2 Results
We ultimately tried all three setups, and the KE results on zsRE dataset with editing number from 0 to 10,00010{,}000 are shown in Tab. 19. For CTFC, as shown in Fig. 6, layers 1 to 5 shares the largest gap between purple bar and green bar, exhibit the largest gap between the purple and green bars, indicating that these layers contribute most significantly to knowledge storage. Thus, editing layers determined by CTFC are layers $1$ to $5$. In addition to Fig. 6, we use heatmap (Fig. 7) visualizations of layer-wise causal effects to analyze how different components of the model contribute to factual knowledge retrieval. These heatmaps guide the selection of editing layers by highlighting consistent and concentrated MLP-specific causal effects in early layers, enabling targeted and effective knowledge editing across architectures.
@@
We observe that for zsRE dataset with 10,00010{,}000 edits, all cases perform badly. As shown in Tab. 19, for the case of CTFC-determined, increasing the editing number from 1,0001{,}000 to 10,00010{,}000 leads to a dramatic drop in ES, falling from $20.17\%$ to just $0.4\%$. For cases using Llama2’s or GPT’s editing layer setting, similar trend also happens: both models’ KE performance drops dramatically to a low level (0.55% for LLama2’s setting, 3.53% for GPT2-XL’s setting) when the editing number rises to 10,00010{,}000. As shown highlighed by orange, DeepSeeK also perform badly on other KE metrics (i.e. PS, NS) when editing number rises to 10,00010{,}000, whereas the smallest PS (0.52%) and NS (2.02%) both appears.
@@ H.1 Analysis
As shown in Tab. 20, LoRA exhibits the smallest average performance decrease ($6.02\%$) compared to DoRA ($6.88\%$) and Full fine-tune ($14.69\%$). In terms of stability, DoRA demonstrates the lowest standard deviation ($7.5$) across models, indicating that the impact of KE on DoRA remains relatively consistent regardless of the base model. Notably, models without fine-tuning behave more erratically: the "No ft" group shows the highest average magnitude of change (Avg_m = $49.46\%$) and standard deviation (Std_m = $169.81$), suggesting high variability. This instability may be attributed to base models’ relatively poor performances on downstream tasks.
Model wise
We found that GPT2-XL demonstrates the most stable performance across different KE methods and datasets among models evaluated. As shown in Fig. 8, GPT2-XL has the smallest varying range from $0.67\%$ to $2.36\%$. In contrast, other models exhibit more variability. For instance, Llama2 experiences a significant fluctuation, with performance ranging from $9.96\%$ to $4.62\%$.
@@
MEMIT generally leads to a smaller reduction in fine-tuning performance compared to AlphaEdit. As illustrated in Fig. 8, AlphaEdit has the largest average decrements of $9.96\%$ which appears on Llama2. A similar pattern is also observed for GPT-J, where AlphaEdit show larger decrements ($6.68\%$ vs $2.52\%$, and $4.57\%$ vs $2.8\%$) across both datasets.
COUNTERFACT dataset tends to cause more severe drops in fine-tuning performance, as evidenced by the top two largest performance declines in Fig. 8 ($9.96\%$, $7.92\%$) occurring in cases involving COUNTERFACT as KE dataset.
Number of edits also introduces a high degree of variability, with different patterns observed across models, as detailed in Tab. 6. For example, GPT-J edited with zsRE using MEMIT shows an improvement in fine-tuning performance from $36.95\%$ to $40.92\%$ as the number of edits increases from $100$ to $1000$, followed by a slight decline to $39.69\%$ when the number of edits reaches 10,00010{,}000. Conversely, GPT-J, edited with zsRE and fine-tuned with the full dataset, shows a reverse trend, with performance dropping from $35.75\%$ to $33.46\%$ and then rising back to $35.5\%$ as the editing number increases. In addition to these curve-like patterns, a common trend observed in some models is a monotonic decrease in performance. For example, in the case of Llama2 edited by AlphaEdit and fine-tuned by DoRA, the performance consistently drops as the number of edits increases, where performance falls from $78.95\%$ ($10^{2}$ edits) to $69.61\%$ ($10^{3}$ edits), and eventually to $63.62\%$ ($10^{4}$ edits).
@@ H.2 Results breakdown
GPT2-XL’s performances on downstream tasks after being edited on zsRE dataset are shown in Tab. 21 (no fine-tune and full fine-tune) and Tab. 22 (LoRA and DoRA). Performances of cases being edited on COUNTERFACT dataset are shown in Tab. 23 (no fine-tune and full fine-tune) and Tab. 24 (LoRA and DoRA)
@@
Llama2’s performances on downstream tasks after being edited on zsRE dataset are shown in Tab. 25 (no fine-tune and full fine-tune) and Tab. 26 (LoRA and DoRA). Performances of cases being edited on COUNTERFACT dataset are shown in Tab. 27 (no fine-tune and full fine-tune) and Tab. 28 (LoRA and DoRA)
@@
Llama3.1
Llama3.1’s performances on downstream tasks after being edited on zsRE dataset are shown in Tab. 29 (LoRA and DoRA). Performances of cases being edited on COUNTERFACT dataset are shown in Tab. 30 (LoRA and DoRA).
zsRE LoRA DoRA MEMIT AlphaEdit MEMIT MEMIT No edit $10^{2}$ $10^{3}$ $10^{4}$ $10^{2}$ $10^{3}$ $10^{4}$ No edit $10^{2}$ $10^{3}$ $10^{4}$ $10^{2}$ $10^{3}$ $10^{4}$ BoolQ 74.35 68.72 59.95 61.17 70.13 61.81 63.12 73.85 67.33 61.86 63.00 69.16 66.50 67.85 PIQA 84.87 78.08 67.29 69.11 81.07 70.53 72.16 88.81 82.39 75.39 77.04 84.37 81.40 82.68 SIQA 78.80 72.49 63.12 64.91 74.86 65.37 67.38 80.27 74.65 68.47 70.01 76.26 73.61 74.73 HellaSwag 91.57 85.16 74.09 76.64 87.00 75.69 78.30 94.20 86.65 79.28 81.05 88.49 85.40 86.69 WinoGrande 84.17 77.44 67.37 69.70 80.36 70.73 72.74 84.67 78.33 71.78 73.42 80.14 77.34 78.54 ARC-e 84.17 77.44 66.69 68.65 79.94 69.55 71.15 90.10 83.19 76.13 77.77 85.59 82.61 83.88 ARC-c 71.20 65.30 56.78 58.77 67.64 59.20 60.78 78.50 72.62 66.41 67.84 74.58 71.98 73.09 openbookqa 80.80 74.40 64.80 66.80 76.80 67.00 69.00 84.80 78.40 72.00 73.60 80.60 77.80 79.00
Table 29: Downstream task performances (%) of Llama3.1 ($M_{\text{ed\_ft}}$) edited on zsRE dataset and then being LoRA or DoRA fine-tuned.
CF1 LoRA DoRA MEMIT AlphaEdit MEMIT MEMIT No edit $10^{2}$ $10^{3}$ $10^{4}$ $10^{2}$ $10^{3}$ $10^{4}$ No edit $10^{2}$ $10^{3}$ $10^{4}$ $10^{2}$ $10^{3}$ $10^{4}$ BoolQ 74.35 68.40 60.19 62.24 70.27 61.33 64.65 73.85 67.20 62.50 65.86 69.42 67.34 68.73 PIQA 84.87 78.08 67.12 70.27 80.63 70.96 75.79 88.81 81.71 76.80 73.54 85.36 82.80 83.65 SIQA 78.80 72.50 63.80 67.43 74.07 66.66 71.11 80.27 73.05 67.21 71.59 76.26 73.21 75.50 HellaSwag 91.57 85.16 74.89 76.64 87.99 77.43 82.71 94.20 86.66 80.59 84.06 88.55 85.89 86.78 WinoGrande 84.17 78.28 68.88 72.02 80.36 71.52 75.54 84.67 77.90 72.45 76.34 80.44 78.03 79.64 ARC-e 84.17 77.44 67.38 69.70 79.96 70.37 75.96 90.10 82.89 77.92 74.60 85.60 83.03 83.89 ARC-c 71.20 65.50 57.64 60.26 67.64 60.19 64.26 78.50 72.00 65.52 69.84 73.79 70.84 72.32 openbookqa 80.80 74.34 65.42 68.39 75.95 67.59 72.15 84.80 78.86 72.56 77.28 80.56 78.14 79.75
Table 30: Downstream task performances (%) of Llama3 ($M_{\text{ed\_ft}}$) edited on COUNTERFACT1 dataset and then being LoRA or DoRA fine-tuned.
@@
GPT-J’s performances on downstream tasks after being edited on zsRE dataset are shown in Tab. 31 (no fine-tune and full fine-tune) and Tab. 32 (LoRA and DoRA). Performances of cases being edited on COUNTERFACT dataset are shown in Tab. 33 (no fine-tune and full fine-tune) and Tab. 34 (LoRA and DoRA)
@@
Appendix I Breakdown of activation-related analysis
The statistical results of layer-wise drift and directional similarity analysis are presented in Tab. 35 and Tab. 36, respectively.
Metrics GPT-2-xl M_ed GPT-2-xl M_ed_tt GPT-J M_ed GPT-J M_ed_tt Llama2 M_ed Llama2 M_ed_tt Max 0.12 0.99 0.99 0.78 1 0.3 Min 0 0.14 0.16 0 0 0.17 Std 0.05 0.36 0.37 0.35 0.54 0.57 Avg 0.04 0.16 0.14 0.21 0.21 0.24
Table 35: Layer-wise drift result breakdown
Metrics GPT2-XL GPT-J Llama2 M_ed - M_ft M_ed_ft - M_ft M_ed_ft - M_ed M_ed - M_ft M_ed_ft - M_ft M_ed_ft - M_ed M_ed - M_ft M_ed_ft - M_ft M_ed_ft - M_ed Max 0.04 0.76 0.15 0.4 0.62 0.63 0.2 0.68 0.24 Min 0 0 0 0 0 0 0 0.44 0 Std. 0 0.51 0.04 0.18 0.52 0.38 0.12 0.6 0.17 Avg. 0.02 0.29 0.05 0.15 0.12 0.2 0.08 0.09 0.07
Table 36: Directional similarity result breakdown
Appendix J Significance Test
We use the t-test to evaluate the statistical significance of different FT methods on KE performance across two dimensions: model-wise and FT method-wise. P-values are computed using paired data, where each pair comprises KE performances of an edited-only model ($M_{ed}$) and its edited-and-fine-tuned counterpart ($M_{ed\_ft}$). In the model-wise dimension, experimental configurations are grouped by model, and for each model, p-values are calculated under various FT methods using paired samples with editing counts ranging from 100 to 10,000. For FT method-wise analysis, configurations are grouped by fine-tuning method, and a p-value is computed for each method group.
Model wise
As shown in Figure 9, GPT‑2 XL exhibits low p‑values under both DoRA and LoRA, indicating that the effect of FT on KE performance is genuine and substantial rather than a product of random variation. For Llama, the impact of full fine‑tuning in reducing KE performance is the most pronounced among all the models. In the case of GPT‑J, although some configurations yield relatively higher p‑values, the majority remain consistently low, pointing to strong effects and suggesting that FT influences GPT‑J in a targeted and effective manner.
Figure 9: Significance test results across models. The format of Y label is KE method-KE dataset, e.g. AE-zsRE means running AlphaEdit on zsRE dataset.
FT method wise
The p-values for DoRA, LoRA, and Full Fine-Tuning are $3.29\times 10^{-5}$, $4.46\times 10^{-5}$, and $0.009$, respectively. These consistently low p-values indicate that the observed performance differences are statistically significant and primarily attributable to the choice of fine-tuning method, rather than intrinsic variation within the editing process itself.
Appendix K Reproducibility related
@@ Appendix I Breakdown of activation-related analysis
    "rewrite_module_tmp":
    "model.layers.{}.mlp.down_proj",
@@
Appendix L Use of Large Language Models